what it the pattern for 120,60,180,90,blank,blank,405
Answer by MathLover1(20850)


120,60,180,90,blank,blank,405
pattern:
120:2=60
60*3=180
180:2=90
90*3=270
270:2=135
135*3=405
so, you have


120,60,180,90,270,135,405